# Conceptualization of this study, Methodology, Software

### v1

@@ 6 Conclusion
By incorporating the Retnet architecture and employing signal embedding for processing 1D EEG signals, this study introduces an innovative methodology aiming to leverage Retnet comprehensively for EEG signal denoising. The integration of Retnet architecture enhances the understanding and processing of temporal information in EEG signals, while the utilization of signal embedding underscores the processing of prolonged temporal information and feature extraction. Experimental results showcase the outstanding denoising performance of our proposed EEGDiR network on EOG and EMG datasets. In comparison to traditional EEG denoising methods, EEGDiR demonstrates notable enhancements in temporal information processing and global modeling.
The global modeling prowess of Retnet, coupled with its favorable handling of temporal information, positions it as an optimal choice for processing EEG signals. The incorporation of signal embedding further refines the representation of EEG signals, preserving context and temporal relationships more effectively. The synergistic application of the Retnet and signal embedding strategy yields a substantial improvement in the denoising performance of the EEGDiR network.

### v2

@@ 6 Conclusion
By incorporating the Retentive Network architecture and employing signal embedding for processing EEG signals, this study introduces an innovative methodology aiming to leverage Retentive Network comprehensively for EEG signal denoising. The integration of Retentive Network architecture enhances the understanding and processing of temporal information in EEG signals, while the utilization of signal embedding underscores the processing of prolonged temporal information and feature extraction. Experimental results showcase the outstanding denoising performance of our proposed EEGDiR network on EOG, EMG and SS2016 EOG datasets. In comparison to traditional EEG denoising methods, EEGDiR demonstrates notable enhancements in temporal information processing and global modeling.
The global modeling prowess of EEGDiR, coupled with its favorable handling of temporal information, positions it as an optimal choice for processing EEG signals. The incorporation of signal embedding further refines the representation of EEG signals, preserving context and temporal relationships more effectively. The synergistic application of the Retentive Network and signal embedding strategy yields a substantial improvement in the denoising performance of the EEGDiR network.
This study holds significant implications as a guide for integrating deep learning into neuroscience, offering valuable insights to enhance the efficacy and application potential of EEG signal processing. By providing an out-of-the-box deep learning dataset, our contribution enables subsequent researchers to expedite EEG signal denoising research by eliminating the need for extensive data preprocessing. This accelerates the development of EEG signal denoising methods.